OpenStudy (anonymous):

A ball is dropped from a height of 10 feet and returns to a height that is one-half of the height from which it fell. How far will the ball have traveled when it hits the ground for the fourth time? A. 20 feet B. 25 feet C. 27.5 feet D. 10 feet

OpenStudy (anonymous):

The distance traveled by the ball is given by the sequence: 10+5+5+2.5+2.5+1.25+1.25= 10+(10+5+2.5). 10+17.5=27.5 feet in total.